OpenStudy (anonymous):

Evaluate f(x) = x 2 + 1 for f(10)

OpenStudy (whpalmer4):

Is that \[f(x) = x^2+1\]To evaluate \(f(10)\) just rewrite the equation putting (10) wherever you see \(x\), then do the arithmetic.

OpenStudy (whpalmer4):

For example\[f(3) = (3)^2+1 = 9+1 = 10\]